# Function theory on the annulus in the dp-norm

### v1

@@ 4. DP-Szegő kernels and normalized DP-Szegő kernels for the tuple $(\lambda_{1},\dots,\lambda_{n})$
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. Every DP-Szegő kernel $[g_{ij}]$ from $\mathcal{G}_{\mathrm{dp}}(\lambda)$ is diagonally congruent to a normalized DP-Szegő kernel.
@@
For any matrix $[g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, we can define a positive definite matrix $[h_{ij}]$ by
@@
On conjugating the inequalities (4.8) by the matrix $C$ we find that $[h_{ij}]$ belongs to $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. ∎
@@
$\begin{bmatrix}(1-\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ (4.25)
@@
$\begin{bmatrix}(1-\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda).$ (4.26)
@@
Definition 4.32.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. Let $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ be a DP Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$. The matrices
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]_{i,j=1}^{n}\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]_{i,j=1}^{n}$ (4.33)
are said to be the defect matrices of $g$.
@@
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0\ \text{and}\ [1-\frac{\delta^{2}}{\overline{z_{i}}z_{j}}g_{ij}]\geq 0$
@@ 5. The DP Pick problem and DP-Szegő kernels
$1-\overline{z_{i}}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(\bar{\lambda}_{i}\lambda_{j}-\delta^{2})b_{ij}$
@@
${\mathcal{C}}=\left\{[(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}]_{i,j=1}^{n}++\left[\left(1-\frac{\delta^{2}}{\bar{\lambda}_{i}\lambda_{j}}\right)b_{ij}\right]_{i,j=1}^{n}:\,[a_{ij}]_{i,j=1}^{n}\geq 0\ \text{and}\ [b_{ij}]_{i,j=1}^{n}\geq 0\right\}\subset H_{n}.$ (5.5)
@@
Note that every $n\times n$ positive semi-definite matrix $[a_{ij}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$. By the positivity of Szegő kernel $\left[\frac{1}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$, the $n\times n$ matrix of the form
$\left[\frac{a_{ij}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$
is also positive semi-definite. In the definition of ${\mathcal{C}}$ (5.5) we may replace $[a_{ij}]_{i,j=1}^{n}$ by $\left[\frac{a_{ij}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$ and $[b_{ij}]_{i,j=1}^{n}$ by the zero matrix, to deduce that $[a_{ij}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$.
By the Hahn-Banach theorem, to show that $[1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$ it suffices to prove that, for every real linear functional ${\mathcal{L}}$ on $H_{n}$, ${\mathcal{L}}\geq 0$ on ${\mathcal{C}}$ implies ${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0$.
Extend ${\mathcal{L}}$ to a complex linear functional $\tilde{\mathcal{L}}$ on $\mathbb{C}^{n\times n}$ by
~$\tilde{\mathcal{L}}(X+{\mathrm{i}}\,Y)={\mathcal{L}}(X)+{\mathrm{i}}\,{\mathcal{L}}(Y)$
@@
$\langle\,c,d\,\rangle_{L}=\tilde{\mathcal{L}}(c\otimes d)$
@@
$\langle\,c,c\,\rangle_{L}=\tilde{\mathcal{L}}(c\otimes c)={\mathcal{L}}(c\otimes c)\geq 0.$
@@
For $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$, we have
$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}$ =~$=\tilde{\mathcal{L}}(x\otimes x)-\tilde{\mathcal{L}}(Tx\otimes Tx)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\lambda_{i}e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\bar{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\bar{\xi}_{i}\bar{\lambda}_{i}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left[(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
$={\mathcal{L}}\left[(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.7)
@@
$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}\geq 0,$
and so $x\in{\mathcal{N}}$ implies that $Tx\in{\mathcal{N}}$. Hence $T$ induces an operator $\tilde{T}$ on $\mathbb{C}^{n}/{\mathcal{N}}$ by
$\tilde{T}(x+{\mathcal{N}})=Tx+{\mathcal{N}},$
and $\|\tilde{T}(x+{\mathcal{N}})\|^{2}\leq\|x+{\mathcal{N}}\|^{2}$ for all $(x+{\mathcal{N}})\in\mathbb{C}^{n}/{\mathcal{N}}$.
Similarly, let us prove that $\|\delta\tilde{T}^{-1}\|\leq 1$. For $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$, we have
$\langle\,x,x\,\rangle_{L}-\langle\,\delta T^{-1}x,\delta T^{-1}x\,\rangle_{L}$ =~$=\tilde{\mathcal{L}}(x\otimes x)-\tilde{\mathcal{L}}(\delta T^{-1}x\otimes\delta T^{-1}x)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\frac{\delta}{\lambda_{j}}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\frac{\delta}{\lambda_{i}}e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\bar{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\frac{\delta}{\lambda_{j}}\bar{\xi}_{i}\frac{\delta}{\overline{\lambda_{i}}}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
$={\mathcal{L}}\left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.8)
Thus
$\langle\,x,x\,\rangle_{L}-\langle\,\delta T^{-1}x,\delta T^{-1}x\,\rangle_{L}\geq 0,$
and so $x\in{\mathcal{N}}$ implies that $\delta T^{-1}x\in{\mathcal{N}}$.
Therefore, $T$ is a Douglas-Paulsen operator. Since the eigenvalues of $T$, $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, $\sigma(T)\subseteq R_{\delta}$, the operator $T\in\mathcal{F}_{\mathrm{dp}}(\delta)$. Therefore, by Proposition 4.9, $[\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}$ is a DP Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$.
By supposition,
$[(1-\bar{z}_{i}z_{j})\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}\geq 0.$
Choose a polynomial $p$ such that $p(\lambda_{i})=z_{i}$, $i=1,\dots,n$. Then ${}p(\tilde{T})e_{i}=z_{i}e_{i}$, $i=1,\dots,n$, and so $\|p(\tilde{T})\|\leq 1$. Choose $c=\begin{bmatrix}1\\ \dots\\ 1\end{bmatrix}\in\mathbb{C}^{n}$. Then
$\langle\,(1-p(\tilde{T})^{*}p(\tilde{T}))c,c\,\rangle_{L}\geq 0,$
@@
${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}*cc^{*})\geq 0,\ \text{and so }\ {\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0,$
@@
${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0.$
Hence $[1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$. ∎
@@
$1-\bar{z}_{i}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\overline{\lambda_{i}}\lambda_{j}})b_{ij}\quad\text{for}\ i,j=1,\dots,n.$ (5.13)
@@
$1+\langle\lambda_{j}x_{j},\lambda_{i}x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle\frac{\delta}{\lambda_{j}}y_{j},\frac{\delta}{\lambda_{i}}y_{i}\rangle_{\mathbb{C}^{r_{2}}}=\bar{z}_{i}z_{j}+\langle x_{j},x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle y_{j},y_{i}\rangle_{\mathbb{C}^{r_{2}}}\ \text{for}\ i,j=1,\dots,n.$
@@
we have $3$ cases:
@@
is not solvable, while, for $\delta=\tfrac{1}{4}$, the Pick interpolation problem is uniquely solvable, and one sees by inspection that the unique solution is the function $f(\lambda)=\lambda$.
@@
A solvable DP-Pick problem on $R_{\delta}$ which is not solvable as a Pick problem on $\mathbb{D}$. Let $\delta\in(0,1)$ and let $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$. We have $10<\delta<\frac{\delta+\sqrt{\delta}}{2}<\sqrt{\delta}<$ 1, so that $\lambda_{1},\lambda_{2}\in R_{\delta}$. Recall from Example 2.9 that the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$. Consider the DP-Pick problem
@@
from which it follows that $D_{1}<0$, and hence $D<0$. Thus the Pick matrix $P$ is not positive, and so, by Pick’s Theorem, the Pick interpolatioin problem $\lambda_{j}\mapsto z_{j},j=1,2$, is not solvable.
@@
Let $\delta\in(0,1)$. The Szegő kernel $[\frac{1}{1-\bar{\mu}\lambda}]$ restricted to $R_{\delta}$ is not a DP Szegő kernel on $R_{\delta}$.

### v2

@@ 4. DP-Szegő kernels and normalized DP-Szegő kernels for the tuple $(\lambda_{1},\dots,\lambda_{n})$
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. Every DP-Szegő kernel $g_{ij}]$ from $\mathcal{G}_{\mathrm{dp}}(\lambda)$ is diagonally congruent to a normalized DP-Szegő kernel.
@@
For any matrix $g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, we can define a positive definite matrix $h_{ij}]$ by
@@
On conjugating the inequalities (4.8) by the matrix $C$ we find that $h_{ij}]$ belongs to $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. ∎
@@
$\begin{bmatrix}(1-\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ (4.24)
@@
$\begin{bmatrix}(1-\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda).$ (4.25)
@@
$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0\ \text{and}\ [1-\frac{\delta^{2}}{\overline{z_{i}}z_{j}}g_{ij}]\geq 0$
@@
Problem 4.31.
Find an effective description of the irreducible DP Szegő kernels on $R_{\delta}$.
@@ 5. The DP Pick problem and DP-Szegő kernels
In this section we shall prove our main theorem, which is a solvability criterion for DP Pick problems in terms of DP-Szegő kernels. We also present some examples which illustrate the relationship between the Pick and DP Pick interpolation problems.
The following notation and terminology will be needed in the proofs.
@@
$1-\overline{z_{i}}z_{j}=(1-\overline{\lambda_{i}}\lambda_{j})a_{ij}+(\overline{\lambda}_{i}\lambda_{j}-\delta^{2})b_{ij}$
@@
${\mathcal{C}}=\left\{[(1-\overline{\lambda_{i}}\lambda_{j})a_{ij}]_{i,j=1}^{n}+\left[\left(1-\frac{\delta^{2}}{\overline{\lambda_{i}}\lambda_{j}}\right)b_{ij}\right]_{i,j=1}^{n}:[a_{ij}]_{i,j=1}^{n}\geq 0\ \text{and}\ [b_{ij}]_{i,j=1}^{n}\geq 0\right\}.$ (5.5)
@@
Note that every $n\times n$ positive semi-definite matrix $a_{ij}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$}. By the positivity of Szegő kernel $\left[\frac{1}{1-\overline{\lambda_{i}}\lambda_{j}}\right]_{i,j=1}^{n}$, the $n\times n$ matrix of the form
$\left[\frac{a_{ij}}{1-\overline{\lambda_{i}}\lambda_{j}}\right]_{i,j=1}^{n}$
is also positive semi-definite. In the definition of $\mathcal{C}$} (5.5) we may replace $a_{ij}]_{i,j=1}^{n}$ by $\left[\frac{a_{ij}}{1-\overline{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$ and $b_{ij}]_{i,j=1}^{n}$ by the zero matrix, to deduce that $a_{ij}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$}.
By the Hahn-Banach theorem, to show that $1-\overline{z_{i}}z_{j}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$} it suffices to prove that, for every real linear functional $\mathcal{L}$} on $H_{n}$, ${\mathcal{L}}\geq 0$ on $\mathcal{C}$} implies ${\mathcal{L}}([1-\overline{z_{i}}z_{j}]_{i,j=1}^{n})\geq 0$.
Extend $\mathcal{L}$} to a complex linear functional $\widetilde{\mathcal{L}}$ on $\mathbb{C}^{n\times n}$ by
$\widetilde{\mathcal{L}}(X+{\mathrm{i}}\,Y)={\mathcal{L}}(X)+{\mathrm{i}}\,{\mathcal{L}}(Y)$
@@
$\langle\,c,d\,\rangle_{L}=\widetilde{\mathcal{L}}(c\otimes d)$
@@
$\langle\,c,c\,\rangle_{L}=\widetilde{\mathcal{L}}(c\otimes c)={\mathcal{L}}(c\otimes c)\geq 0.$
@@
Let us construct an operator $\widetilde{T}$ on $\mathbb{C}^{n}/{\mathcal{N}}$ such that $\|\widetilde{T}\|\leq 1$ and $\|\delta\widetilde{T}^{-1}\|\leq 1$. For $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$, we have
$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}$ $=\widetilde{\mathcal{L}}(x\otimes x)-\widetilde{\mathcal{L}}(Tx\otimes Tx)$
$=\widetilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\widetilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\lambda_{i}e_{i}\right)$
$=\widetilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\overline{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\widetilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\overline{\xi}_{i}\overline{\lambda}_{i}e_{j}\,\otimes e_{i}\right)$
$=\widetilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}(1-\overline{\lambda}_{i}\lambda_{j})\overline{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
$=\widetilde{\mathcal{L}}\left[(1-\overline{\lambda}_{i}\lambda_{j})\overline{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
$={\mathcal{L}}\left[(1-\overline{\lambda}_{i}\lambda_{j})\overline{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.7)
@@
$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}\geq 0,$ (5.8)
and so $x\in{\mathcal{N}}$ implies that $Tx\in{\mathcal{N}}$. Hence $T$ induces an operator $\widetilde{T}$ on $\mathbb{C}^{n}/{\mathcal{N}}$ by
$\widetilde{T}(x+{\mathcal{N}})=Tx+{\mathcal{N}},$
and $\|\widetilde{T}(x+{\mathcal{N}})\|^{2}\leq\|x+{\mathcal{N}}\|^{2}$ for all $(x+{\mathcal{N}})\in\mathbb{C}^{n}/{\mathcal{N}}$, which implies that
$\|\widetilde{T}\|\leq 1.$ (5.9)
Notice that $\sigma(T)=\{\lambda_{1},\dots,\lambda_{n}\}\subset R_{\delta}$ and so $T$ is invertible. Moreover
$\delta T^{-1}e_{j}=\frac{\delta}{\lambda_{j}}e_{j}\ \text{for}\ j=1,\dots,n,$
and so, in the chain of equations leading to equation (5), we may replace $T$ by $\delta T^{-1}$ and $\lambda_{j}$ by $\frac{\delta}{\lambda_{j}}$ to deduce that, for $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$,
$\langle x,x\rangle_{L}-\langle\delta T^{-1}x,\delta T^{-1}x\rangle_{L}=\mathcal{L}\left[\left(1-\tfrac{\delta}{\overline{\lambda_{i}}}\tfrac{\delta}{\lambda_{j}}\right)\overline{\xi_{i}}\xi_{j}\right]_{j,i=1}^{n}.$
Clearly $\left[\left(1-\tfrac{\delta}{\overline{\lambda_{i}}}\tfrac{\delta}{\lambda_{j}}\right)\overline{\xi_{i}}\xi_{j}\right]_{j,i=1}^{n}\in\mathcal{C}$ (take $a_{ij}=0,b_{ij}=\overline{\xi_{i}}\xi_{j}$ in the defining expression (5.5)), and so, since $\mathcal{L}\geq 0$ on $\mathcal{C}$, we have
$\langle x,x\rangle_{L}-\langle\delta T^{-1}x,\delta T^{-1}x\rangle_{L}\geq 0.$ (5.10)
Thus $x\in{\mathcal{N}}$ implies that $\delta T^{-1}x\in{\mathcal{N}}$, and therefore $\delta T^{-1}$ induces an operator $\widetilde{(\delta T^{-1})}$ on $\mathbb{C}^{n}/\mathcal{N}$ by
$\widetilde{(\delta T^{-1})}(x+\mathcal{N})=\delta T^{-1}x+\mathcal{N},$
and in the light of inequality (5.10),
$\|\widetilde{(\delta T^{-1})}\|\leq 1.$ (5.11)
We have, for any $x\in\mathbb{C}^{n}$,
$\widetilde{T}\widetilde{(\delta T^{-1})}(x+\mathcal{N})=\widetilde{T}(\delta T^{-1}x+\mathcal{N})=T\delta T^{-1}x+\mathcal{N}=\delta(x+\mathcal{N}),$
and so $\widetilde{(\delta T^{-1})}=\delta(\widetilde{T})^{-1}$. Hence, by the inequality (5.11),
$\|\delta(\widetilde{T})^{-1}\|=\|\widetilde{(\delta T^{-1})}\|\leq 1.$
Therefore, $\tilde{T}$ is a Douglas-Paulsen operator. Since the eigenvalues of $T$, which are $\lambda_{1},\dots,\lambda_{n}$, belong to $R_{\delta}$, $\sigma(T)\subseteq R_{\delta}$, and so the operator $T$ belongs to $\mathcal{F}_{\mathrm{dp}}(\delta,\lambda)$. Therefore, by Proposition 4.9, $\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
Let $g_{ij}=\langle\,e_{j},e_{i}\,\rangle_{L}$ for $i,j=1,\dots,n$. By supposition (5.4),
$(1-\overline{z}_{i}z_{j})\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}\geq 0.$
Choose a polynomial $p$ such that $p(\lambda_{i})=z_{i}$, $i=1,\dots,n$. Then $p(\widetilde{T})e_{i}=z_{i}e_{i}$, $i=1,\dots,n$. Observe that
$\left[\langle(1-p(\tilde{T})^{*}p(\tilde{T}))e_{j},e_{i}\rangle\right]$ $=\left[\langle e_{j},e_{i}\rangle-\langle p(\tilde{T})e_{j},p(\tilde{T})e_{i}\rangle\right]$
$=\left[\langle e_{j},e_{i}\rangle-\langle z_{j}e_{j},z_{i}e_{i}\rangle\right]$
$=\left[(1-\overline{z}_{i}z_{j})g_{ij}\right]\geq 0.$
Therefore, $\|p(\widetilde{T})\|\leq 1$. Choose $c=\begin{bmatrix}1\\ \dots\\ 1\end{bmatrix}\in\mathbb{C}^{n}$. Then
$\langle\,(1-p(\widetilde{T})^{*}p(\widetilde{T}))c,c\,\rangle_{L}\geq 0,$
@@
${\mathcal{L}}([1-\overline{z}_{i}z_{j}]_{i,j=1}^{n}*cc^{*})\geq 0,\ \text{and so }\ {\mathcal{L}}([1-\overline{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0,$
@@
${\mathcal{L}}([1-\overline{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0.$
Hence $1-\overline{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$}. ∎
We show in the next theorem that, as in the classical Pick theorem, if a DP Pick problem is solvable then it is solvable by a rational function in $\mathcal{S}_{\mathrm{dp}}$.
@@
$1-\overline{z}_{i}z_{j}=(1-\overline{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\overline{\lambda_{i}}\lambda_{j}})b_{ij}\quad\text{for}\ i,j=1,\dots,n.$ (5.16)
@@
$1+\langle\lambda_{j}x_{j},\lambda_{i}x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle\frac{\delta}{\lambda_{j}}y_{j},\frac{\delta}{\lambda_{i}}y_{i}\rangle_{\mathbb{C}^{r_{2}}}=\overline{z}_{i}z_{j}+\langle x_{j},x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle y_{j},y_{i}\rangle_{\mathbb{C}^{r_{2}}}\ \text{for}\ i,j=1,\dots,n.$
@@
As the dp norm and sup norm are different, the converse statement to Remark 5.21 is false, as one would expect. The following two examples provide concrete instances of this fact.
@@
solvability depends on the value of $\delta$. There are $3$ cases:
@@
is not solvable, while, for $\delta=\tfrac{1}{4}$, the Pick interpolation problem is uniquely solvable, and one sees by inspection that the unique solution is the function $f(\lambda)=\lambda$. ∎
@@
Another solvable DP Pick data-set which is not a solvable Pick data-set on $\mathbb{D}$. Let $\delta\in(0,1)$ and let $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$. We have $0<\delta<\frac{\delta+\sqrt{\delta}}{2}<\sqrt{\delta}<1$, so that $\lambda_{1},\lambda_{2}\in R_{\delta}$. Recall from Example 2.9 that the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$. Consider the DP-Pick problem
@@
from which it follows that $D_{1}<0$, and hence $D<0$. Thus the Pick matrix $P$ is not positive, and so, by Pick’s Theorem, the Pick interpolation problem $\lambda_{j}\mapsto z_{j},j=1,2$, is not solvable.
Since the Pick interpolation problem on $\mathbb{D}$ and the DP Pick problem on $R_{\delta}$ are so closely related, it is natural to ask whether the Szegő kernel on $\mathbb{D}$, when retricted to $R_{\delta}$, is a DP-Szegő kernel. We can use Example 5.25 to answer this question in the negative.
@@
Let $\delta\in(0,1)$. The Szegő kernel $\frac{1}{1-\overline{\mu}\lambda}]$ restricted to $R_{\delta}$ is not a DP Szegő kernel on $R_{\delta}$.